# Polynomial-time isomorphism test for $k$-generated extensions of abelian groups

### v1

@@ 5.1 Finding an isomorphism
Additive generators of $K$ can be computed in polynomial time by solving a system of linear Diophantine equations, see Proposition 2.2. By Theorem 1.6, we can compute multiplicative generators of the unit group
@@
Since $g_{1}A,\dots,g_{k}A$ generate $G/A$, we can express each element $g\in G$ in the form $g=w(g_{1},\dots,g_{k})x$, where $w$ is some word and $x\in A$. Moreoever, $w$ and $x$ can be found in polynomial time (although they are not unique). Given this decomposition of $g$ we can compute its image under $\Phi$ as $\Phi(g)=w(g_{1}^{0},\dots,g_{k}^{0})\phi(x)$. It follows that we can compute $\Phi$ in polynomial time, and it is easy to check if $\Phi$ is an isomorphism by bruteforce. ∎
@@ 5.2 Computing the coset of isomorphisms
As in Step 4 of Section 5.1, we can compute permutation generators of $K$ in polynomial time by solving a system of linear Diophantine equations and applying Theorem 1.6. Let $C$ be the pointwise stabilizer of $u_{1},\dots,u_{m}$ in $K$, note that it can be computed in polynomial time.
@@ 7 Acknowledgements
The author expresses his gratitude to A.V. Vasil’ev for fruitful discussions on the topic of this paper.

### v2

@@ 4 Computing unit groups of finite rings
We will apply Theorem 1.6 to the case when the ring in question acts on an abelian group by endomorphisms.
Let $A$ be a finite abelian group given by its Cayley table. Let $\operatorname{End}(A)$ be the ring of endomorphisms of $A$. We can encode every element $f\in\operatorname{End}(A)$ as a tuple $(f(a))_{a\in A}$ of $|A|$ elements of $A$, so the additive group $\operatorname{End}(A)^{+}$ can be viewed as a subgroup of the direct power $A^{|A|}$. The size of each element in this representation is polynomial in $|A|$, and we can compute element sums and products (compositions) in polynomial time. Note that in general the order of $\operatorname{End}(A)$ is not polynomial in $|A|$, but we have a bound $|\operatorname{End}(A)|\leq|A|^{\log|A|}$.
Let $\operatorname{End}(A)^{+}$ denote the additive group of $\operatorname{End}(A)$, and let $e_{1},\dots,e_{t}$ be cyclic generators of $A$. For each $a\in A$ set $N(a)=\{b\in A\mid b^{|a|}=1\}$. For $i=1,\dots,t$ and $a\in N(e_{i})$ we can define an endomorphism $f_{i,a}$ of $A$ by $f_{i,a}(e_{1}^{x_{1}}\cdots e_{t}^{x_{t}})=a^{x_{i}}$, where $x_{1},\dots x_{t}\in\mathbb{Z}$. Note that $f_{i,a}$, $i=1,\dots,t$, $a\in N(e_{i})$, generate $\operatorname{End}(A)^{+}$ additively. One can compute the cyclic generators of $\operatorname{End}(A)^{+}$ and express any given element in terms of these generators in time polynomial in $|A|$.
Let $R$ be a unitary subring of $\operatorname{End}(A)$ given by its additive generators. Note that $R^{\times}$ acts on $A$ by automorphisms, in particular, every element of $R^{\times}$ is a permutation on $A$. Hence we can specify $R^{\times}$ as a permutation group in terms of its permutation generators; their number is polynomial in $|A|$. All prime divisors of $|\operatorname{End}(A)|$ also divide $|A|$, hence the largest prime divisor of $|R|$ is bounded by $|A|$.
Theorem 1.6 implies the following.
Corollary 4.3.
@@ 5.1 Finding an isomorphism
Additive generators of $K$ can be computed in polynomial time by solving a system of linear Diophantine equations, see Proposition 2.2. By Corollary 4.3, we can compute multiplicative generators of the unit group
@@
Since $g_{1}A,\dots,g_{k}A$ generate $G/A$, we can express each element $g\in G$ in the form $g=w(g_{1},\dots,g_{k})x$, where $w$ is some word and $x\in A$. Moreover, $w$ and $x$ can be found in polynomial time (although they are not unique). Given this decomposition of $g$ we can compute its image under $\Phi$ as $\Phi(g)=w(g_{1}^{0},\dots,g_{k}^{0})\phi(x)$. It follows that we can compute $\Phi$ in polynomial time, and it is easy to check if $\Phi$ is an isomorphism by bruteforce. ∎
@@ 5.2 Computing the coset of isomorphisms
As in Step 4 of Section 5.1, we can compute permutation generators of $K$ in polynomial time by solving a system of linear Diophantine equations and applying Corollary 4.3. Let $C$ be the pointwise stabilizer of $u_{1},\dots,u_{m}$ in $K$, note that it can be computed in polynomial time.
@@ 7 Acknowledgements
The author expresses his gratitude to A.V. Vasil’ev and I.N. Ponomarenko for fruitful discussions on the topic of this paper and many useful suggestions.